OpenStudy (anonymous):

Can some solve this problem for me : 3^4x-7 = 4^2x+3 I have no idea how to do it.

OpenStudy (amistre64):

been solved , by many people, thru many days

OpenStudy (amistre64):

those exponents? if so, group them as such

OpenStudy (amistre64):

3^a = 4^b a = b ln(4)/ln(3) might be useful

OpenStudy (amistre64):

4x - 7 = (2x+3)c 4x - 7 = 2xc+3c 4x -2xc = 7 +3c x(4 -2c) = 7 +3c x = (7 +3c)/(4 -2c)